Let x ∈ R such that

(2021 −2020/n)≤ x ≤( 2021 +2020/n)

for all positive integer n ≥ 10. Prove that x = 2021.


1
Expert's answer
2021-01-25T01:45:13-0500

"\lim_{n->\infin}(2021-\frac{2020}n{})\le x\le\lim_{n->\infin}(2021+\frac{2020}{n})"

"2021\le x\le 2021"

"x=2021"